owns 9,000 acres of timberland purchased in E11-19 (L04) (Depletion Computations-Timber) Stanislaw Timber Company 2006 at a cost of $1,400 per acre. At the time of purchase, the land without the timber was valued at $400 per acre. In 2007 Stanislaw built fire lanes and roads, with a life of 30 years, at a cost of $84,000. Every year, Stanislaw sprays to prevent disease at a cost of S3,000 per year and spends $7,000 to maintain the fire lanes and roads. During 2008, Stanislaw selectively logged an sold 700,000 board feet of timber, of the estimated 3,500,000 board feet. In 2009, Stanislaw planted new seedlings to replace th trees cut at a cost of $100,000. Instructions (a) Determine the depreciation expense and the cost of timber sold related to depletion for 2008.
Answer #1
a) $84,000/ 30 = $2,800 depreciation expense per year
===============================================
b) per acre - land value = Cost of Timber Sold
1400 - 400 = 1000
===============================================
Cost of Timber Sold x acres = Timber Value
1000 x 9000 = 9000000
===============================================
Timber value/ Est. Board Feet   x sold board feet =land value
9000000/3500000 x 700000                 =1800000
===============================================
Timber Value -Land Value = Cost of Timber Sold
9000000 1800000 = 7200000
===============================================
Cost of Timber Sold +Replacement Cost
7200000 +100000 =7300000
===============================================
Est. Board Feet 5,000,000
                       1.46 x
Timber Value 900000
===============================================
Depletion 1314000